OpenStudy (anonymous):

Let f(x)=10x+(arctanx)^2+2. If g(x) = f-1(x) <- inverse. If g is the inverse function of f, then find the value of g'(2) so far I have got the derivative of g(x) f'(y) = 10y'+2(arctany)y'/1+y^2 and I can't figure out what to do with the y^2

OpenStudy (dumbcow):

\[g'(2) = \frac{1}{f'(g(2))}\] \[f'(x) = 10 + \frac{2 \tan^{-1} x}{1+x^{2}}\] g(2) is the solution to f(x) = 2 \[10x + (\tan^{-1} x)^{2} +2 = 2\] \[10x +(\tan^{-1} x)^{2} = 0\] x = 0 http://www.wolframalpha.com/input/?i=10x+%2B+arctan%28x%29%5E2+%3D+0 \[f'(0) = 10 + 2\tan^{-1} 0 / 1+0= 10\] \[g'(2) = \frac{1}{10}\]

OpenStudy (anonymous):

thank you so much!

OpenStudy (dumbcow):

no problem... sorry i kinda got lost trying to follow your work, i like to just keep everything in terms of x

OpenStudy (anonymous):

It was my fault should've made it more clearer :(

OpenStudy (dumbcow):

you took the derivative correctly though just looks like you used implicit differentiation ?

OpenStudy (anonymous):

yes I did implicit differentiation, but that just got me more confused. Your method was much easier to understand the question :)

OpenStudy (dumbcow):

:)